# A Survey of NL2SQL with Large Language Models: Where are we, and where are we going?

### v4

@@ VI-A SQL Correction Strategies
The sql generated by nl2sql models may contain syntax errors or unnecessary keywords like DESC, DISTINCT, and Aggregate functions. DIN-SQL [5] introduces a self-correction module that operates in a zero-shot setting, where the model receives only the faulty sql and attempts to correct it. Two prompts are used: a general prompt for CodeX, which directly asks for error identification and correction, and a mild prompt for GPT-4, which seeks potential issues without presuming errors. To handle errors in predicate predictions, such as incorrect columns or values, ZeroNL2SQL [44] employs a multi-level matching approach. This method incrementally expands matching across columns, tables, and databases, allowing matched values to be returned to the LLMs to generate sql queries consistent with the database content.
While these methods focus on fixing syntax errors, they often overlook semantic errors, such as incorrect table joins, misaligned conditions, or inaccurate aggregations, which are essential for improving accuracy.
@@ VI-B Output Consistency
C3-SQL [62] introduces a Consistency Output (CO) component to reduce LLM output randomness and improve zero-shot nl2sql performance. CO samples multiple reasoning paths, executes them, filters errors, and applies voting to select the most consistent sql result. DAIL-SQL [6] integrates self-consistency, achieving a $0.4\%$ performance improvement over configurations without it. To reduce LLM randomness, Fin-SQL [50] generates $n$ candidate sql queries in parallel, clusters them based on keyword consistency, and selects a query from the largest cluster. The self-consistency method enhances LLM output diversity by raising the temperature and then using majority voting to select the final result.
However, recent studies [117] suggest that a single model may not provide sufficient diversity. To address this, PET-SQL [57] introduces a cross-consistency strategy, in which multiple LLMs generate sql at lower temperatures and vote based on execution results.
While these methods improve consistency across multiple runs with single or multiple LLMs, reducing model randomness and enhancing accuracy, they significantly increase inference cost and time.
@@ VI-C Execution-Guided Strategies
In nl2sql tasks, the execution result of an sql query provides critical feedback on nl2sql translation accuracy. For example, errors or NULL values in execution results can signal potential issues with the sql query.
@@ VI-D N-best Rerankers Strategies
In nl2sql tasks, especially in cross-domain scenarios, generated sql queries can vary subtly in structure and semantics. N-best reranking reorders the top-$N$ model outputs, often leveraging a larger model or additional knowledge sources. For example, fine-tuning a BERT-based reranker, as demonstrated by Bertrand-dr on the Spider dataset [118], has effectively improved the performance of several nl2sql models.
@@
VII NL2SQL Benchmarks
In this section, we will first elaborate on the different types of nl2sql datasets, highlighting their characteristics, as shown in Figure 9 (Section VII-A). We will then perform an in-depth analysis of existing nl2sql datasets (Section VII-B).
VII-A An Overview of NL2SQL Benchmarks
With advancements in nl2sql, various datasets have emerged to address the evolving challenges, as shown in Figure 9. These range from single-domain databases with simple queries to cross-domain, multi-turn, multilingual, and domain-specific scenarios, reflecting the progress and the emergence of new challenges for nl2sql solutions.
Figure 9: Timeline for NL2SQL Benchmarks.
TABLE II: Statistics of NL2SQL Benchmarks.
@@ VII-A An Overview of NL2SQL Benchmarks
Single-Domain NL2SQL Datasets. Early nl2sql datasets focused on specific domains with relatively simple sql queries, such as ATIS [121] for flight information and GeoQuery [122] for U.S. geographical facts. Recently, larger single-domain datasets [130, 132, 141, 146, 50, 147] have been introduced, featuring more complex databases and sql queries tailored to specific scenarios. This shift reflects an increased emphasis on assessing nl2sql systems’ performance and practical utility within particular domains.
@@
Cross-Domain NL2SQL Datasets. After the development of early single-domain datasets, the nl2sql field shifted toward cross-domain datasets to test systems’ generalization across diverse sql queries and databases. WikiSQL [107] was the first cross-domain dataset, drawing tables from Wikipedia across various domains. Subsequently, Spider [45] was introduced, containing more complex relational databases with multiple tables. Recently, the BIRD [52] has further advanced complexity by including sql functions and operations absent in Spider, providing an even greater challenge for nl2sql.
@@
Multi-Turn NL2SQL Datasets. With advancements in nl2sql, multi-turn datasets have been developed to support interactive dialogues. SParC [127] is a cross-domain and multi-turn dataset with about $4.3K$ nl questions, totaling over $12K$ (nl, sql) pairs, each nl questions requiring contextual understanding across turns. CoSQL [128], collected using a Wizard-of-Oz setup, includes over $30K$ turns and introduces additional challenges like unanswerable questions, further testing context comprehension. CHASE [136], the first multi-turn Chinese dataset, reduces context-independent questions and simple sql queries, advancing dialogue-centered nl2sql research.
@@
NL2SQL Datasets with Robustness Testing. In real-world applications, nl2sql systems must handle diverse user groups and databases, emphasizing robustness. Spider-Syn [137] simulates user unfamiliarity with schemas by using synonyms in nl questions, while Dr.Spider [144] applies 17 types of perturbations to databases, nl questions, and sql queries for comprehensive robustness evaluation.
@@
NL2SQL Datasets with SQL Efficiency Testing. Real-world databases often hold vast amounts of data, and multiple sql queries can answer a single user question, each with varying sql execution efficiency. BIRD [52] introduces a metric for evaluating SQL execution efficiency called the Valid Efficiency Score (VES), which will be further discussed in Section VIII.
@@
Knowledge-Augmented NL2SQL Datasets. Domain-specific knowledge is essential for nl2sql systems to perform well in real-world applications. KaggleDBQA [140] includes database documents, such as column and table descriptions. Similarly, Spider-DK [138] expands the Spider development set by adding five types of domain knowledge to nl questions, testing systems’ ability to understand and use this information.
@@
NL2SQL Datasets with Ambiguous Questions. In real-world nl2sql tasks, ambiguities often arise, such as semantic ambiguities in nl and overlapping database schemas, making ambiguity-focused evaluation increasingly important. AmbiQT [145] is the first dataset designed to assess ambiguity coverage, comprising four ambiguity types. Each nl question maps to two valid sql queries, reflecting specific ambiguities.
@@
Synthetic NL2SQL Datasets. MIMICSQL [130] employs a template-based approach to generate initial template questions and corresponding sql queries, though manual refinement is required to make questions more natural. ScienceBenchmark [146] also uses templates for initial sql generation but leverages GPT-3 for sql-to-nl translation.
VII-B In-depth Analysis of Existing NL2SQL Datasets
@@
To analyze and compare nl2sql datasets complexity, we use the NL2SQL360 [46] system for statistical evaluation, as shown in Table II, We measure the Redundancy, including the number of nl questions, sql queries, and their ratio. DB Complexity covers the total databases, total tables, average tables per database, average columns per table, and average records per database. Query Complexity measures the average number of tables, SELECT keywords, aggregate functions, scalar functions, and mathematical computations in each sql query. For datasets without public dev/test splits, such as CHASE [136], only statistics for public splits are reported. For datasets without publicly available data, like knowSQL [99], values in Table II are marked with “–”.
@@
Discussion. Despite the increasing number of datasets proposed by the nl2sql community, a gap in sql complexity remains compared to real-world scenarios. Current datasets typically feature fewer SELECT keywords, indicating a lack of nested queries and complex set operations. Additionally, challenges involving Scalar Functions and Mathematical Computations require further focus. We encourage the community to propose datasets addressing these complexities.
@@ VIII Evaluation and Error Analysis
In this section, we introduce key evaluation metrics for nl2sql solutions (Section VIII-A), review toolkits for low-cost and comprehensive evaluation (Section VIII-B), and provide an error taxonomy for analyzing sql errors in the nl2sql process (Section VIII-C).
@@ VIII-A Evaluation Metrics
Evaluation metrics are crucial for assessing the performance of nl2sql solution. We define $N$ as the dataset size, $Q_{i}$ as the nl question of the $i$-th example, $V_{i}$ as the execution result set of the ground-truth sql query $Y_{i}$ and $\hat{V}_{i}$ as the execution result set of the sql query $\hat{Y}_{i}$ generated by the nl2sql solution.
Execution Accuracy (EX) [45]. This metric evaluates the performance of the nl2sql system by comparing whether the execution result sets of the ground-truth sql queries and the predicted sql queries are identical. It can be computed by: $EX=\frac{\sum_{i=1}^{N}\mathbbm{1}(V_{i}=\hat{V}_{i})}{N},$ where $\mathbbm{1}(\cdot)$ is an indicator function that equals 1 if the condition inside is satisfied, and 0 otherwise. Note that false negatives could occur because different sql queries corresponding to semantically different nl queries may produce identical execution result sets.
@@
Component-Match Accuracy (CM) [45]. This metric evaluates the detailed performance of the nl2sql system by measuring the exact matching of different sql components such as SELECT, WHERE and others between the ground-truth sql query and the predicted sql query. For a specific sql component $C$. The computation can be formalized as follows:
$CM^{C}=\frac{\sum_{i=1}^{N}\mathbbm{1}(Y_{i}^{C}=\hat{Y}_{i}^{C})}{N},$
@@
$EM=\frac{\sum_{i=1}^{N}\mathbbm{1}(\bigwedge_{C_{k}\in\mathbb{C}}Y_{i}^{C_{k}}=\hat{Y}_{i}^{C_{k}})}{N}.$
@@
$\small VES=\frac{\sum_{i=1}^{N}\mathbbm{1}(V_{i}=\hat{V}_{i})\cdot\mathbf{R}(Y_{i},\hat{Y_{i}})}{N},~\mathbf{R}(Y_{i},\hat{Y_{i}})=\sqrt{\frac{\mathbf{E}(Y_{i})}{\mathbf{E}(\hat{Y}_{i})}},$
@@
Query Variance Testing (QVT) [46]. This metric measures the robustness of an nl2sql system in handling variations in nl queries. For a given sql query $Y_{i}$, there are often multiple corresponding nl queries, represented as pairs {($Q_{1}$, $Y_{i}$), ($Q_{2}$, $Y_{i}$), …, ($Q_{m}$, $Y_{i}$)}. The QVT metric is calculated as:
@@
VIII-B NL2SQL Evaluation Toolkits
@@
Recent nl2sql solutions have achieved remarkable performance on various nl2sql benchmarks. However, in real-world applications, variations in nl query styles, database schemas, and sql query characteristics across domains make it difficult to fully assess system robustness using standard benchmark metrics alone. To address this, recent toolkits [142, 46] have been developed to provide a more comprehensive evaluation of nl2sql systems in practical scenarios.
@@
MT-TEQL [142] is a unified framework for evaluating the performance of nl2sql systems in handling real-world variations in nl queries and database schemas. It is based on a metamorphic testing approach, implementing semantic-preserving transformations of nl queries and database schemas to generate their variants without manual efforts automatically. It includes four types of transformations for nl queries (e.g. Prefix Insertion) and eight types of transformations for database schemas: (e.g. Table Shuffle).
@@
NL2SQL360 [46] is a multi-angle evaluation framework offering fine-grained assessments of nl2sql systems across diverse scenarios (Figure LABEL:fig:overview(c)). Unlike MT-TEQL, it emphasizes varied sql query characteristics in different applications, such as aggregate functions, nested queries, or top-$k$ queries typical of the Business Intelligence scenario. Comprising six core components, i.e., Dataset, Model Zoo, Metrics, Dataset Filter, Evaluator, and Analysis, NL2SQL360 provides a unified, model-agnostic interface for systematic evaluations. Users can apply both public and private datasets, customize metrics for specific scenarios, and analyze performance on subsets with scenario-specific sql characteristics, offering valuable insights into nl2sql system effectiveness across applications.
VIII-C A Taxonomy for NL2SQL Errors Analysis
@@
Error analysis involves examining model errors to identify limitations and guide corrective actions for improved performance. In this section, we first review the existing nl2sql error taxonomy. We then propose design principles and introduce a two-level nl2sql errors taxonomy.
Existing Taxonomies for NL2SQL Errors Analysis. Recent nl2sql research [150, 151, 54, 5, 152] has increasingly incorporated error analysis, proposing various error taxonomies.
Ning et al. [151] introduced a detailed error taxonomy based on two dimensions: (1) Syntactic dimension identifies specific sql parts where errors occur, organized by keywords such as WHERE and JOIN. (2) Semantic dimension indicates misinterpretations of the natural language description, such as errors in understanding table names. SQL-PaLM [152] categorizes errors into five types: (1) Schema Linking, where the model fails to select the relevant tables required for nl; (2) Misunderstanding Database Content, where the model is not able to accurately interpret the data within the tables; (3) Misunderstanding Knowledge Evidence, where the model fails to use or ignores human-annotated evidence; (4) Reasoning, which includes an inadequate understanding of the question, leading to queries that lack the necessary reasoning steps to generate correct results; and (5) Syntax-Related Errors, where the generated sql contains syntax errors.
@@
Figure 10: A Data-Driven Roadmap and a Decision Flow for Recommending NL2SQL modules.
@@
Our Taxonomy for NL2SQL Errors Analysis. Current error taxonomies in nl2sql are often specific to particular datasets, limiting their general applicability. To address these issues, a standardized and effective taxonomy is essential. We propose the following principles to guide the development of an nl2sql error taxonomy:
@@
  • Comprehensiveness: It should cover all possible error types in the nl2sql translation process.
@@
  • Extensibility: The taxonomy should be adaptable to include emerging error types as nl2sql evolves.
@@
Following these principles, we developed a two-level nl2sql error analysis taxonomy:
@@
A Case Study of Error Analysis. We collected and classified errors from DIN-SQL [5] on the Spider [45] using our proposed taxonomy. As shown in Figure LABEL:fig:overview(d), the results suggest that this taxonomy is effective. However, we acknowledge that achieving a comprehensive nl2sql error taxonomy requires ongoing refinement. We invite the research community to contribute to the continuous improvement of this taxonomy.
IX Practical Guidance for NL2SQL Solutions
In this section, we provide practical guidance for developing nl2sql solutions, considering key factors and scenarios.
IX-A Data-Driven Roadmap of Optimizing LLMs for NL2SQL
In Figure 10(a), we outline a strategic roadmap designed to optimize LLMs for nl2sql task, based on data privacy and data volume. Data privacy affects the choice of open-source and closed-source LLMs, while data volume affects the strategies for optimization for training and inference.

### v5

@@ VI-A SQL Correction Strategies
The sql generated by Text-to-SQL models may contain syntax errors. DIN-SQL [5] introduces a self-correction module that operates in a zero-shot setting, where the model receives only the faulty sql and attempts to correct it. Two prompts are used: a general prompt for CodeX, which directly asks for error identification and correction, and a mild prompt for GPT-4, which seeks potential issues without presuming errors. To handle errors in predicate predictions, such as incorrect columns or values, ZeroNL2SQL [44] employs a multi-level matching approach. This method incrementally expands matching across columns, tables, and databases, allowing matched values to be returned to the LLMs to generate sql queries consistent with the database content.
While these methods focus on fixing syntax errors, they often overlook semantic errors [33], such as incorrect table joins, misaligned conditions, or inaccurate aggregations, which are essential for improving accuracy.
@@ VI-B Output Consistency
DAIL-SQL [6] integrates self-consistency, achieving a $0.4\%$ performance improvement over configurations without it. To reduce LLM randomness, Fin-SQL [50] generates $n$ candidate sql queries in parallel, clusters them based on keyword consistency, and selects a query from the largest cluster. The self-consistency strategy enhances LLM output diversity by increasing the temperature and selecting the final result through majority voting. However, recent studies [118] indicate that relying on a single model may still yield limited diversity. To overcome this limitation, PET-SQL [62] introduces a cross-consistency strategy, in which multiple LLMs generate sql at lower temperatures and vote based on execution results.
While these methods improve accuracy by enforcing consistency across multiple executions, they significantly increase inference cost and time.
@@ VI-C Execution-Guided Strategies
In Text-to-SQL tasks, the execution result of an sql query provides critical feedback on Text-to-SQL translation accuracy. For example, errors or NULL values in execution results can signal potential issues with the sql query.
@@ VI-D N-best Rerankers Strategies
In Text-to-SQL tasks, especially in cross-domain scenarios, generated sql queries can vary subtly in structure and semantics. N-best reranking reorders the top-$N$ model outputs, often leveraging a larger model or additional knowledge sources. For example, fine-tuning a BERT-based reranker, as demonstrated by Bertrand-dr on the Spider dataset [119], has effectively improved the performance of several Text-to-SQL models.
@@
VII Text-to-SQL Benchmarks
In this section, we will first elaborate on the different types of Text-to-SQL datasets, highlighting their characteristics, as shown in Figure 9 (Section VII-A). We will then perform an in-depth analysis of existing datasets (Section VII-B).
VII-A An Overview of Text-to-SQL Benchmarks
With advancements in Text-to-SQL, various datasets have emerged to address the evolving challenges, as shown in Figure 9. These range from single-domain databases with simple queries to cross-domain, multi-turn, multilingual, and domain-specific scenarios, reflecting the progress and the emergence of new challenges for Text-to-SQL solutions.
TABLE II: Statistics of Text-to-SQL Benchmarks.
@@ VII-A An Overview of Text-to-SQL Benchmarks
Single-Domain Text-to-SQL Datasets. Early Text-to-SQL datasets focused on specific domains with relatively simple sql queries, such as ATIS [122] for flight information and GeoQuery [123] for U.S. geographical facts. Recently, larger single-domain datasets [131, 133, 142, 147, 50, 148] have been introduced, featuring more complex databases and sql queries tailored to specific scenarios. This shift reflects an increased emphasis on assessing Text-to-SQL systems’ performance and practical utility within particular domains.
@@
Cross-Domain Text-to-SQL Datasets. After the development of early single-domain datasets, the Text-to-SQL field shifted toward cross-domain datasets to test systems’ generalization across diverse sql queries and databases. WikiSQL [110] was the first cross-domain dataset, drawing tables from Wikipedia across various domains. Subsequently, Spider [45] was introduced, containing more complex relational databases with multiple tables. Recently, the BIRD [52] has further advanced complexity by including sql functions and operations absent in Spider, providing a greater challenge for Text-to-SQL.
@@
Multi-Turn Text-to-SQL Datasets. With advancements in Text-to-SQL, multi-turn datasets have been developed to support interactive dialogues. SParC [128] is a cross-domain and multi-turn dataset with about $4.3K$ nl questions, totaling over $12K$ (nl, sql) pairs, each nl questions requiring contextual understanding across turns. CoSQL [129], collected using a Wizard-of-Oz setup, includes over $30K$ turns and introduces additional challenges like unanswerable questions, further testing context comprehension.
@@
Text-to-SQL Datasets with Robustness Testing. In real-world applications, Text-to-SQL systems must handle diverse user groups and databases, emphasizing robustness. Spider-Syn [138] simulates user unfamiliarity with schemas by using synonyms in nl questions, while Dr.Spider [145] applies 17 types of perturbations to databases, nl questions, and sql queries for comprehensive robustness evaluation.
@@
Text-to-SQL Datasets with SQL Efficiency Testing. Real-world databases often hold vast amounts of data, and a single nl may correspond to multiple sql queries with different execution efficiencies. BIRD [52] introduces a metric for evaluating SQL execution efficiency called the Valid Efficiency Score (VES), which will be further discussed in Section VIII.
@@
Knowledge-Augmented Text-to-SQL Datasets. Domain-specific knowledge is essential for Text-to-SQL systems to perform well in real-world applications. KaggleDBQA [141] includes database documents, such as column and table descriptions. Similarly, Spider-DK [139] expands the Spider development set by adding five types of domain knowledge to nl questions, testing systems’ ability to use this information.
@@
Text-to-SQL Datasets with Ambiguous Questions. In real-world Text-to-SQL tasks, ambiguities often arise, such as semantic ambiguities in nl and overlapping database schemas, making ambiguity-focused evaluation increasingly important. AmbiQT [146] is the first dataset designed to assess ambiguity coverage, comprising four ambiguity types. Each nl question maps to two valid sql queries, reflecting specific ambiguities.
@@
Synthetic Text-to-SQL Datasets. MIMICSQL [131] employs a template-based approach to generate initial template questions and corresponding sql queries, though manual refinement is required to make questions more natural. ScienceBenchmark [147] also uses templates for initial sql generation but leverages GPT-3 for sql-to-nl translation.
VII-B In-depth Analysis of Existing Text-to-SQL Datasets
@@
To analyze and compare Text-to-SQL datasets complexity, we use the NL2SQL360 [46] system for statistical evaluation, as shown in Table II, We measure the Redundancy, including the number of nl questions, sql queries, and their ratio. DB Complexity covers the total databases, total tables, average tables per database, average columns per table, and average records per database. Query Complexity measures the average number of tables, SELECT keywords, aggregate functions, scalar functions, and mathematical computations in each sql query. For datasets without public dev/test splits, such as CHASE [137], only statistics for public splits are reported. For datasets without publicly available data, like knowSQL [102], values in Table II are marked with “–”.
@@
Discussion. Despite the increasing number of datasets proposed by the Text-to-SQL community, a gap in sql complexity remains compared to real-world scenarios. Current datasets typically feature fewer SELECT keywords, indicating a lack of nested queries and complex set operations. Additionally, challenges involving Scalar Functions and Mathematical Computations require further focus. We encourage the community to propose datasets addressing these complexities.
@@ VIII Evaluation and Error Analysis
In this section, we introduce key evaluation metrics for Text-to-SQL solutions (Section VIII-A), review toolkits for low-cost and comprehensive evaluation (Section VIII-B), and provide an error taxonomy for analyzing sql errors in the Text-to-SQL process (Section VIII-C).
@@ VIII-A Evaluation Metrics
Evaluation metrics are crucial for measuring Text-to-SQL performance. We define $N$ as the dataset size, $Q_{i}$ as the nl question of the $i$-th example, $V_{i}$ as the execution result set of the ground-truth sql query $Y_{i}$ and $\hat{V}_{i}$ as the execution result set of the sql query $\hat{Y}_{i}$ generated by the Text-to-SQL solution.
Execution Accuracy (EX) [45]. This metric evaluates the performance of the Text-to-SQL system by comparing whether the execution result sets of the ground-truth sql queries and the predicted sql queries are identical. It can be computed by: $EX=\frac{\sum_{i=1}^{N}\mathbbm{1}(V_{i}=\hat{V}_{i})}{N},$ where $\mathbbm{1}(\cdot)$ is an indicator function that equals 1 if the condition inside is satisfied, and 0 otherwise. Note that false negatives could occur because different sql queries corresponding to semantically different nl queries may produce identical execution result sets.
@@
Component-Match Accuracy (CM) [45]. This metric evaluates the detailed performance of the Text-to-SQL system by measuring the exact matching of different sql components such as SELECT, WHERE and others between the ground-truth sql query and the predicted sql query. For a specific sql component $C$. The computation can be formalized as follows:
$CM^{C}=\frac{\sum_{i=1}^{N}\mathbbm{1}(Y_{i}^{C}=\hat{Y}_{i}^{C})}{N},\vskip-5.0pt$
@@
$EM=\frac{\sum_{i=1}^{N}\mathbbm{1}(\bigwedge_{C_{k}\in\mathbb{C}}Y_{i}^{C_{k}}=\hat{Y}_{i}^{C_{k}})}{N}.\vskip-5.0pt$
@@
$\small VES=\frac{\sum_{i=1}^{N}\mathbbm{1}(V_{i}=\hat{V}_{i})\cdot\mathbf{R}(Y_{i},\hat{Y_{i}})}{N},~\mathbf{R}(Y_{i},\hat{Y_{i}})=\sqrt{\frac{\mathbf{E}(Y_{i})}{\mathbf{E}(\hat{Y}_{i})}},\vskip-4.62497pt$
@@
Query Variance Testing (QVT) [46]. This metric measures the robustness of an Text-to-SQL system in handling variations in nl queries. For a given sql query $Y_{i}$, there are often multiple corresponding nl queries, represented as pairs {($Q_{1}$, $Y_{i}$), ($Q_{2}$, $Y_{i}$), …, ($Q_{m}$, $Y_{i}$)}. The QVT metric is calculated as:
@@
VIII-B Text-to-SQL Evaluation Toolkits
@@
Recent Text-to-SQL solutions have achieved remarkable performance on various Text-to-SQL benchmarks. However, in real-world applications, variations in nl query styles, database schemas, and sql query characteristics across domains make it difficult to fully assess system robustness using standard benchmark metrics alone. To address this, recent toolkits [143, 46] have been developed to provide a more comprehensive evaluation of Text-to-SQL systems in practical scenarios.
@@
MT-TEQL [143] is a unified framework for evaluating the performance of Text-to-SQL systems in handling real-world variations in nl queries and database schemas. It is based on a metamorphic testing approach, implementing semantic-preserving transformations of nl queries and database schemas to generate their variants without manual efforts automatically. It includes four types of transformations for nl queries (e.g. Prefix Insertion) and eight types of transformations for database schemas: (e.g. Table Shuffle).
@@
NL2SQL360 [46] is a multi-angle evaluation framework offering fine-grained assessments of Text-to-SQL systems across diverse scenarios (Figure LABEL:fig:overview(c)). Unlike MT-TEQL, it emphasizes varied sql query characteristics in different applications, such as aggregate functions, nested queries, or top-$k$ queries typical of the Business Intelligence scenario. Comprising six core components, Dataset, Model Zoo, Metrics, Dataset Filter, Evaluator, and Analysis, NL2SQL360 provides a unified, model-agnostic interface for systematic evaluations. Users can apply both public and private datasets, customize metrics for specific scenarios, and analyze performance on subsets with scenario-specific sql characteristics, offering valuable insights into Text-to-SQL system effectiveness across applications.
VIII-C A Taxonomy for Text-to-SQL Errors Analysis
@@
Error analysis involves examining model errors to identify limitations and guide corrective actions for improved performance. In this section, we first review the existing Text-to-SQL error taxonomy. We then propose design principles and introduce a two-level Text-to-SQL errors taxonomy.
Existing Taxonomies for Text-to-SQL Errors Analysis. Recent Text-to-SQL research [33, 151, 152, 59, 5, 153] has increasingly incorporated error analysis, proposing various error taxonomies. Ning et al. [152] introduced a detailed error taxonomy based on two dimensions: (1) Syntactic dimension identifies specific sql parts where errors occur, organized by keywords such as WHERE and JOIN. (2) Semantic dimension indicates misinterpretations of the natural language description, such as errors in understanding table names. SQL-PaLM [153] categorizes errors into five types: (1) Schema Linking, irrelevant or missing table/column selection; (2) Database Content, misinterpreting data values; (3) Knowledge Evidence, failing to utilize external hints; (4) Reasoning, lacking intermediate logical steps; and (5) Syntax, invalid SQL format. NL2SQL-BUGs [33] focuses on the analysis of semantic errors, organizing them into 9 main categories and 31 subcategories. It further proposes a new benchmark for evaluating models’ error detection capabilities, advancing automated error analysis in the Text-to-SQL.
@@
Figure 10: A Data-Driven Roadmap and a Decision Flow for Recommending Text-to-SQL modules.
@@
Taxonomy Principles for Text-to-SQL Errors Analysis. Current error taxonomies in Text-to-SQL are often specific to particular datasets, limiting their general applicability. To address these issues, a standardized and effective taxonomy is essential. We propose the following principles [154] to guide the development of an Text-to-SQL error taxonomy:
@@
  • Comprehensiveness: The taxonomy should cover all possible error types in the Text-to-SQL translation process.
@@
  • Extensibility: The taxonomy should be adaptable to include emerging error types as Text-to-SQL evolves.
@@
Our Taxonomy for Text-to-SQL Errors Analysis. Following these principles, we developed a two-level Text-to-SQL error analysis taxonomy:
@@
Discuss the Application of the Two-level Error Taxonomy. We collected and classified errors from DIN-SQL [5] on the Spider [45] using our proposed taxonomy. As shown in Figure LABEL:fig:overview(d), only 1.8% of the errors fall into the Others category, suggesting our taxonomy is practical and effective.
Nonetheless, we recognize that developing a complete and universally applicable Text-to-SQL error taxonomy is inherently iterative. We encourage continued efforts from the community to refine and expand this taxonomy over time.
IX Practical Guidance for Text-to-SQL
In this section, we provide practical guidance for developing Text-to-SQL solutions, considering key factors and scenarios.
IX-A Data-Driven Roadmap for Text-to-SQL
In Figure 10(a), we outline a strategic roadmap designed to optimize LLMs for Text-to-SQL task, based on data privacy and data volume. Data privacy affects the choice of open-source and closed-source LLMs, while data volume affects the strategies for optimization for training and inference.